# Test-Time Anchoring for Discrete Diffusion Posterior Sampling

### v1

@@ Appendix A Additional Theoretical Results
From the proof of Theorem A.1, the conditional likelihood satisfies
@@
$\mathrm{KL}\!\Big(q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})\,\|\,p_{\theta}(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})\Big)=\mathrm{KL}\!\Big(q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})\,\|\,q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\theta}(Z_{t(i)}))\Big).$
Following the argument of Theorem A.1, we distinguish two cases for $Z_{t(i)}$. For $Z_{t(i)}={\mathbf{m}}$, the KL-divergence evaluates to
@@
Substituting the above expression in the conditional likelihood (A), we get
@@
$p_{\varphi}(Z_{s(i)}|Z_{t(i)},{\mathbf{y}})=q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\,q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)})).$
@@
$\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log\frac{p_{\theta}(Z_{s(i)}|Z_{t(i)})}{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\,q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))}\Bigg]$
$=\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log\frac{p_{\theta}(Z_{s(i)}|Z_{t(i)})}{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))}-\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\Bigg]$
$=\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log\frac{p_{\theta}(Z_{s(i)}|Z_{t(i)})}{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))}\Bigg]-\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\Bigg].$
@@
$\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log\frac{p_{\theta}(Z_{s(i)}|Z_{t(i)})}{p_{\varphi}(Z_{s(i)}|Z_{t(i)},{\mathbf{y}})}\Bigg]$ (13)
$=\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log\frac{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\theta}(Z_{t(i)}))}{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))}\Bigg]-\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\Bigg].$ (14)
Similarly to the proof of Theorem A.1, we consider two cases to compute the first expectation. Case I: $Z_{t(i)}={\mathbf{m}}$. Using the masked diffusion inference posterior derived from (3), when $Z_{t(i)}={\mathbf{m}}$ we have
@@
which simplifies (13) as follows:
@@
$=\frac{\alpha_{s(i)}-\alpha_{t(i)}}{1-\alpha_{t(i)}}\,\log\frac{\langle{\mathbf{x}}_{\theta}(Z_{t(i)}),{\mathbf{x}}\rangle}{\langle{\mathbf{x}}_{\varphi}(Z_{t(i)}),{\mathbf{x}}\rangle}\bm{1}_{\{Z_{t(i)}={\mathbf{m}}\}}\,-\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\Bigg].$ (15)
Substituting (A) into (A), the conditional likelihood can be bounded as
@@
$\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\Big],$ (16)
where in the last step we use the fact that $q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))$ is independent of $Z_{s(i)}$ given $Z_{t(i)}$.
Treatment of boundary conditions. In masked diffusion the terminal state $Z_{1}$ is absorbing (all-mask), hence $q(Z_{1}|{\mathbf{x}})=p_{\varphi}(Z_{1}|{\mathbf{y}})$ and the boundary KL-divergence vanishes. Thus (A) becomes
$-\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})$ $\leq\mathbb{E}_{q(Z_{0:1}|{\mathbf{x}})}\!\big[-\log p_{\varphi}({\mathbf{x}}|Z_{0},{\mathbf{y}})\big]$
@@
$\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big].$ (17)
Since $\mathbf{x}$} is a deterministic function of $Z_{0}$, the reconstruction term in (A) simplifies to:
@@
$\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big].$ (18)
Finally, the decoder is same for unconditional generation and posterior sampling, i.e., $p_{\varphi}({\mathbf{x}}|Z_{0})=p_{\theta}({\mathbf{x}}|Z_{0})$. Substituting this property in (A) yields
$-\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})$ $\leq\mathbb{E}_{q(Z_{0:1}|{\mathbf{x}})}\!\big[-\log p_{\theta}({\mathbf{x}}|Z_{0})\big]$
@@
$\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big]\coloneqq{\mathcal{L}}_{\mathrm{APS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ (19)
Since $\mathrm{CE}\!\big({\mathbf{x}},{\mathbf{x}}_{\theta}(Z_{t(i)})\big)=-\log\langle\rangle$, the first two terms in (A) equals to the standard NELBO (5) used to train the masked diffusion model. Therefore, we have
${\mathcal{L}}_{\mathrm{APS}}({\mathbf{x}},{\mathbf{y}};\varphi)={\mathcal{L}}_{\mathrm{NELBO}}({\mathbf{x}};\theta)$ $+\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Bigg[\frac{\alpha_{s(i)}-\alpha_{t(i)}}{1-\alpha_{t(i)}}\,\log\frac{\langle{\mathbf{x}}_{\theta}(Z_{t(i)}),{\mathbf{x}}\rangle}{\langle{\mathbf{x}}_{\varphi}(Z_{t(i)}),{\mathbf{x}}\rangle}\,\bm{1}_{\{Z_{t(i)}={\mathbf{m}}\}}\Bigg]$
$\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))\big],$
completing the statement of the theorem. ∎
Implications. Theorem A.2 establishes a principled upper bound on the negative log-posterior likelihood when posterior sampling is performed without additional training.
@@
  • Measurement consistency. The final summation enforces alignment with the measurement likelihood $q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}))$, ensuring that the sampler accounts for observations at each diffusion step.
@@
Appendix B Additional Experiments
This section provides supplementary details and evaluations of our APS method. We first describe implementation details for both inverse problems and stylization (§B.1), followed by algorithmic analysis including pseudocode and hyperparameter studies (§B.2). We then examine the impact of design choices in our ablations (§B.3), evaluate using the same prior (§B.4), and discuss computational complexity (§B.5). Next, we outline the compared baselines (§B.6) and summarize benchmarks and metrics (§B.7). Finally, we present additional results (§B.8) and limitations (§B.9).
B.1 Implementation Details
B.1.1 Inverse Problems
@@
At each of the $15$ reverse steps, we perform an inner optimization loop to ensure measurement consistency. This loop consists of $100$ optimization steps using the Adam optimizer with a learning rate of $\eta=1.0$. The total loss function is a weighted sum of three components: (1) a reconstruction loss, which could be L1 or L2 norm; we choose L1 norm $\|{\mathbf{y}}-{\mathcal{A}}(\hat{{\mathbf{x}}})\|_{1}$ because it is known to generate sharper quality (Van Den Oord et al., 2017; Yu et al., 2024), (2) a VGG perceptual loss with a coefficient of $\lambda_{\mathrm{p}}=10^{-3}$, and (3) a prior preservation loss ${\mathcal{L}}_{\mathrm{prior}}({\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)}),{\mathbf{x}}_{\theta}({\mathbf{z}}_{t(i)}))=-\frac{1}{L}\sum_{l=1}^{L}\log\langle{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)}),{\mathbf{x}}_{\theta}({\mathbf{z}}_{t(i)})\rangle\bm{1}_{\{{\mathbf{z}}^{l}_{t(i)}={\mathbf{m}}\}}$ with coefficient $\lambda_{\mathrm{pp}}=10^{-4}$. To reduce the number of hyper-parameters introduced by our Algorithm 1 and ensure fast convergence, we make the following practical implementation choices:
  • initialize ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)})$ with ${\mathbf{x}}_{\theta}({\mathbf{z}}_{t(i)})$: Lines 6 and 9,
  • choose a small decaying learning rate $\eta=1.0$ in Adam: Line 15,
  • optimize for a fewer iterations $T=15$ and $M=100$: Lines 5 and 8.
This has an equivalent effect of prior preservation. We note that the optimization loop (Lines 8-16) computes gradients with respect to the logits $\varphi$. This is significantly cheaper than the expensive backpropagation steps through denoising network (here, 8B parameters for MMaDA) used in prior continuous diffusion approaches (Rout et al., 2023; 2024; Chung et al., 2024). Thus, we set $\lambda_{\mathrm{pp}}=0$ by default. Alternatively, one could initialize $\varphi$ with all zeros or randomly and set $M$ very large until convergence while enforcing the prior preservation loss ($\lambda_{\mathrm{pp}}\neq 0$) in Line 14.
@@
B.1.2 Reference-based Stylization

### v2

@@ Appendix B Additional Theoretical Results
From the proof of Theorem B.1, the conditional likelihood satisfies
$-\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})$ $\leq\mathbb{E}_{q(Z_{0}|{\mathbf{x}})}\!\big[-\log p_{\varphi}({\mathbf{x}}|Z_{0},{\mathbf{y}})\big]+\mathbb{E}_{q(Z_{1}|{\mathbf{x}})}\!\Big[\log\frac{q(Z_{1}|{\mathbf{x}})}{p_{\varphi}(Z_{1}|{\mathbf{y}})}\Big]$
@@
$\mathrm{KL}\!\Big(q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})\,\|\,p_{\theta}(Z_{s(i)}|Z_{t(i)})\Big)=\mathrm{KL}\!\Big(q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})\,\|\,q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\theta}(Z_{t(i)}))\Big).$
Following the argument of Theorem B.1, we distinguish two cases for $Z_{t(i)}$. For $Z_{t(i)}={\mathbf{m}}$, the KL-divergence evaluates to
@@
$\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log\frac{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}{p_{\varphi}(Z_{s(i)}|Z_{t(i)},{\mathbf{y}})}\Bigg]$ (13)
@@
Substituting the above expression in the conditional likelihood (B), we get
$-\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})$ $\leq\mathbb{E}_{q(Z_{0}|{\mathbf{x}})}\!\big[-\log p_{\varphi}({\mathbf{x}}|Z_{0},{\mathbf{y}})\big]+\mathbb{E}_{q(Z_{1}|{\mathbf{x}})}\!\Big[\log\frac{q(Z_{1}|{\mathbf{x}})}{p_{\varphi}(Z_{1}|{\mathbf{y}})}\Big]$
@@
$p_{\varphi}(Z_{s(i)}|Z_{t(i)},{\mathbf{y}})$ $\;{=}\;\frac{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\,q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))}{{\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})}},\quad\text{where}$
${\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})}$ ${\;=\;\sum_{{\mathbf{z}}_{s}\in{\mathcal{V}}}q\!\left({\mathbf{z}}_{s}|Z_{t(i)},\mathbf{x}_{\varphi}(Z_{t(i)})\right)\;q\!\left(\mathbf{y}|{\mathbf{x}}_{\varphi}(Z_{t(i)};{\mathbf{z}}_{s})\right).}$
@@
$\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log\frac{p_{\theta}(Z_{s(i)}|Z_{t(i)}){\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})}}{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))\,q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))}\Bigg]$
$=\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log\frac{p_{\theta}(Z_{s(i)}|Z_{t(i)})}{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))}-\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}})){+\log\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})}\Bigg]$
$=\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log\frac{p_{\theta}(Z_{s(i)}|Z_{t(i)})}{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))}\Bigg]-\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))\Bigg]$
$\hskip 204.85974pt{+\log\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})}.$
@@
$=\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log\frac{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\theta}(Z_{t(i)}))}{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}}_{\varphi}(Z_{t(i)}))}\Bigg]-\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))\Bigg]$
$\hskip 204.85974pt{+\log\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})}.$ (15)
Similarly to the proof of Theorem B.1, we consider two cases to compute the first expectation.
Case I: $Z_{t(i)}={\mathbf{m}}$. Using the masked diffusion inference posterior derived from (3), when $Z_{t(i)}={\mathbf{m}}$ we have
@@
which simplifies (B) as follows:
@@
$=\frac{\alpha_{s(i)}-\alpha_{t(i)}}{1-\alpha_{t(i)}}\,\log\frac{\langle{\mathbf{x}}_{\theta}(Z_{t(i)}),{\mathbf{x}}\rangle}{\langle{\mathbf{x}}_{\varphi}(Z_{t(i)}),{\mathbf{x}}\rangle}\bm{1}_{\{Z_{t(i)}={\mathbf{m}}\}}\,-\mathbb{E}_{q(Z_{s(i)}|Z_{t(i)},{\mathbf{x}})}\!\Bigg[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))\Bigg]$
$\hskip 193.47882pt{+\log\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})}.$ (16)
Substituting (B) into (B), the conditional likelihood can be bounded as
@@
$\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}{,Z_{s(i)}}|{\mathbf{x}})}\Big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))\Big]$
$\quad{\;+\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Big[\log\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})\Big]}.$ (17)
Treatment of boundary conditions. In masked diffusion the terminal state $Z_{1}$ is absorbing (all-mask), hence $q(Z_{1}|{\mathbf{x}})=p_{\varphi}(Z_{1}|{\mathbf{y}})$ and the boundary KL-divergence vanishes. Thus (B) becomes
$-\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})$ $\leq\mathbb{E}_{q(Z_{0}|{\mathbf{x}})}\!\big[-\log p_{\varphi}({\mathbf{x}}|Z_{0},{\mathbf{y}})\big]$
@@
$\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}{,Z_{s(i)}}|{\mathbf{x}})}\Big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))\Big]$
$\quad{\;+\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Big[\log\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})\Big]}.$ (18)
Since $\mathbf{x}$} is a deterministic function of $Z_{0}$, the reconstruction term in (B) simplifies to:
@@
$\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}{,Z_{s(i)}}|{\mathbf{x}})}\Big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))\Big]$
$\quad{\;+\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Big[\log\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})\Big]}.$ (19)
Finally, the decoder is same for unconditional generation and posterior sampling, i.e., $p_{\varphi}({\mathbf{x}}|Z_{0})=p_{\theta}({\mathbf{x}}|Z_{0})$. Substituting this property in (B) yields
$-\log p_{\varphi}({\mathbf{x}}|{\mathbf{y}})$ $\leq\mathbb{E}_{q(Z_{0}|{\mathbf{x}})}\!\big[-\log p_{\theta}({\mathbf{x}}|Z_{0})\big]$
@@
$\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}{,Z_{s(i)}}|{\mathbf{x}})}\Big[\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))\Big]$
$\quad{\;+\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Big[\log\mathcal{Z}_{\varphi}(Z_{t(i)},\mathbf{y})\Big]}\coloneqq{\mathcal{L}}_{\mathrm{APS}}({\mathbf{x}},{\mathbf{y}};\varphi)$ (20)
Note that $\mathrm{CE}\!\big({\mathbf{x}},{\mathbf{x}}_{\theta}(Z_{t(i)})\big)=-\log\langle{\mathbf{x}},{\mathbf{x}}_{\theta}(Z_{t(i)}\rangle$. Generalizing this to $L>1$, the first two terms in (B) equals the standard NELBO (5) used to train the masked diffusion model. Therefore, we have
${\mathcal{L}}_{\mathrm{APS}}({\mathbf{x}},{\mathbf{y}};\varphi)={\mathcal{L}}_{\mathrm{NELBO}}({\mathbf{x}};\theta)$ $+\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Bigg[\frac{\alpha_{s(i)}-\alpha_{t(i)}}{1-\alpha_{t(i)}}\,\sum_{l=1}^{L}\log\frac{\langle{\mathbf{x}}^{l}_{\theta}(Z_{t(i)}),{\mathbf{x}}^{l}\rangle}{\langle{\mathbf{x}}^{l}_{\varphi}(Z_{t(i)}),{\mathbf{x}}^{l}\rangle}\,\bm{1}_{\{Z^{l}_{t(i)}={\mathbf{m}}\}}\Bigg]$
$)]\displaystyle\quad-\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}{,Z_{s(i)}}|{\mathbf{x}})}\Big[\sum_{l=1}^{L}\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z^{l}_{s(i)}}$$)\Big$]
$\quad{\;+\sum_{i=1}^{T}\mathbb{E}_{q(Z_{t(i)}|{\mathbf{x}})}\Big[\sum_{l=1}^{L}\log\mathcal{Z}^{l}_{\varphi}(Z_{t(i)},\mathbf{y})\Big]},$
which upon regrouping completes the statement of the theorem. ∎
Implications. Theorem B.2 establishes a principled upper bound on the negative log-posterior likelihood when posterior sampling is performed without additional training.
@@
  • Measurement consistency. The final summation enforces alignment with the measurement likelihood $q({\mathbf{y}}|{\mathbf{x}}_{\varphi}(Z_{t(i)}{;Z_{s(i)}}))$, ensuring that the sampler accounts for observations at each diffusion step.
@@
Appendix C Additional Experiments
This section provides supplementary details and evaluations of our APS method. We first describe implementation details for both inverse problems and stylization (§C.1), followed by algorithmic analysis including pseudocode and hyperparameter studies (§C.2). We then examine the impact of design choices in our ablations (§C.3), evaluate using the same prior (§C.4), and discuss computational complexity (§C.5). Next, we outline the compared baselines (§C.7) and summarize benchmarks and metrics (§C.8). Finally, we present additional results (§C.9), limitations (§C.10), and conclude with reproducibility (§C.11).
C.1 Implementation Details
C.1.1 Inverse Problems
@@
At each of the $15$ reverse steps, we perform an inner optimization loop to ensure measurement consistency. This loop consists of $100$ optimization steps using the Adam optimizer with a learning rate of $1.0$. The total loss function is a weighted sum of two components: (1) a reconstruction loss, which could be L1 or L2 norm, but we choose L1 norm: $\|{\mathbf{y}}-{\mathcal{A}}(\hat{{\mathbf{x}}})\|_{1}$ as it is known to generate sharper quality), and (2) a VGG perceptual loss with a coefficient of $10^{-3}$.
@@
C.1.2 Reference-based Stylization